Canavan disease is an autosomal recessive disease that causes progressive damage to nerve cells in the brain. Symptoms such as mental retardation, loss of motor skills and abnormal muscle tone progress rapidly beginning in infancy. The disease is fatal with no cure, and death usually occurs before age 4. This disease is caused by mutations in the gene for aspartoacylase on chromosome 17. In the Ashkenazi Jewish population, 1 in 10,000 persons are born with the disease. Assuming that this population has random mating and is in equilibrium, what is the spontaneous mutation rate from normal aspartoacylase to the disease-causing allele? a. 1 x 10-8 b. 2 x 10-8 c. 1 x 10-6 d. 2 x 10-6 e. 1 x 10-4 The answer is e but i would like to understand why and the concept behind it. so if the concept can be explained that would be great thanks. and the step by step math too please.
Answer #1

Canavan disease is autosomal reessive disease

  • Considering the diseased allele = a, normal allele =A

  • Forward mutation rate (A-a) = µ = 5 x 10-5

  • Backward mutation rate = (a-A) = ν = 3 x 10-3

  • p = allele frequency of q = allele frequency of a

Considering equilibrium (no other change)

q= µ / (µ + ν)

Given

Number of diseased individuals == 1in10000 = aa

Frequency of diseased individual = 1/10000 = 10-4

In the given case, there is no reverse mutation.

Thus,

Mutation rate = 1 x 10-4